R^{5}RS Errata

Richard Kelsey September 29, 1999

Several inconsistencies in R^{5}RS have been pointed out since its initial publication. This document addresses the most important of them.

Semantics of call-with-current-continuation

The definition of call-with-current-continuation in section 7.2 is incorrect because it is incompatible with dynamic-wind. As shown in section 4 of [1], however, this incorrect semantics is adequate to define the shift and reset operators, which can then be used to define the correct semantics of both

dynamic-wind and call-with-current-continuation.

Redefinition of keywords

As stated in section 1.2 keywords are not reserved. The grammar specified in section 7.1 is thus somewhat misleading because a program may redefine any or all syntactic keywords. In particular, it is possible to use any identifier as a variable.

Port?

Section 3.2 includes port? in the list of predicates defining disjoint types; there is no such procedure. Input-port? and output-port? should be included in this list of predicates. An object may satisfy both input-port? and output-port?, but cannot satisfy any other pair of predicates in this list.

Begin at top level

As with define, there are two different semantics for begin depending on context. When used as an expression a begin form contains expressions and is library syntax as described in section 4.2.3. At top level a begin form contains commands and definitions and its semantics are as described in section 5.1: a top-level begin is equivalent to the sequence of commands and definitions it contains. A top-level begin is syntax, not library syntax.

A begin at top level that contains only expressions is syntactically ambiguous, but the two semantics are equivalent in that particular case.

List-ref

The description of (list-ref list k) in section 3.2.2 should say

It is an error if list has k or fewer elements. and not

It is an error if list has fewer than k elements.

Reference

[1] Andrzej Filinski. Representing Monads. In Proceedings of the 1994 ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages, pages 446–457.